# Nearly Minimax Optimal Regret for Multinomial Logistic Bandit

### v6

@@ I.2 Proofs of Lemmas for Proposition 1
$\mu(x^{\top}\mathbf{w}_{V})-\mu(\hat{x}^{\top}\mathbf{w}_{V})$ $=\int_{v=1}^{1}\dot{\mu}\left(x^{\top}\mathbf{w}_{V}+v(\hat{x}-x)^{\top}\mathbf{w}_{V}\right)\left(x-\hat{x}\right)^{\top}\mathbf{w}_{V}$
@@
where in the first equality, $\bar{x}$ is the convex combination of $x$ and $\hat{x}$, the first inequality holds by Lemma I.3, the second inequality holds by the bounded assumption (Assumption 1), and the last inequality holds by the definition of $\delta=d/4-|\tilde{U}_{t}\cap V|$. ∎
@@ I.3 Technical Lemmas for Proposition 1
Lemma I.5 (Lemma 11 of Goyal and Perivier [24]).
@@
$\sum_{t=1}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})p_{t}(0|S_{t},\mathbf{w}^{\star})\leq\sum_{t=1}^{T}\kappa_{t}^{\star}+\textup{{Reg}}_{T}(\mathbf{w}^{\star}).$
@@
Then, we can rewrite the right-hand side of (J.1) as follows:
@@
where $\bar{\mathbf{w}}_{t}$ is the convex combination of $\mathbf{w}^{\star}$ and $\mathbf{w}_{t+1}$. The first term of (J.2) can be bound by
@@
Moreover, the second term of (J.2) can be bounded as follows:
@@
where the first inequality holds by Lemma 1, and the last inequality holds by Lemma E.2. Combining (J.3) and (J.4), we can bound the term $\sum_{t\in J_{1}}g_{t}(S_{t};\mathbf{w}^{\star})$ as follows:
@@
Now, by plugging (J.5) and (J.6) into (J.1), and using the fact that $\beta_{T}(\delta)=\tilde{\mathcal{O}}(\sqrt{d})$, we obtain that
@@
Furthermore, by applying the same analysis as in the proof of Theorem 2 in order to bound the second-order term in the Taylor expansion (term (B) in (E.1)), we get
@@
For each instance, we sample the true parameter $\mathbf{w}^{\star}$ from a uniform distribution in $[-1/\sqrt{d},1/\sqrt{d}]^{d}$. For the context features $x_{ti}$, we sample each $x_{ti}$ independently and identically distributed (i.i.d.) from a multivariate Gaussian distribution $\mathcal{N}(\mathbf{0}_{d},\mathbf{I}_{d})$ and clip it to range $[-1/\sqrt{d},1/\sqrt{d}]^{d}$. Therefore, we ensure that $\|\mathbf{w}^{\star}\|_{2}\leq 1$ and $\|x_{ti}\|_{2}\leq 1$, satisfying Assumption 1. For each experimental configuration, we conducted $20$ independent runs for each instance and reported the average cumulative regret (Figure 1) and runtime per round (Figure K.1) for each algorithm. The error bars in Figure 1 and K.2 represent the standard deviations (1-sigma error). We have omitted the error bars in Figure K.1 because they are minimal.
@@
It appears that the authors may have intended to derive this equation using a first-order exact Taylor expansion. However, in MNL bandits, this equation generally does not hold. Consider a counterexample where $x_{ti}=0$, $x_{tj}\neq 0$ for $j\neq i$, and $\theta^{\star}\neq\theta_{t}$. Then, the left-hand side of (L.1) is equals to 00 (since $x_{ti}=0$), but the right-hand side of (L.1) is not 00, because the denominators of each $\mu_{i}(\mathbf{X}_{S_{t}}^{\top}\theta^{\star})$ and $\mu_{i}(\mathbf{X}_{S_{t}}^{\top}\theta_{t})$ differ. This equation only holds in special cases, such as when $K=1$, which corresponds to the logistic bandit case. Equation (16) in Agrawal et al. [4] serves as the foundation for the entire proof in their paper. Consequently, all subsequent results derived from it are also incorrect.

### v9

@@ I.2 Proofs of Lemmas for Proposition 1
$\mu(x^{\top}\mathbf{w}_{V})-\mu(\hat{x}^{\top}\mathbf{w}_{V})$ $=\int_{v=0}^{1}\dot{\mu}\left(x^{\top}\mathbf{w}_{V}+v(\hat{x}-x)^{\top}\mathbf{w}_{V}\right)\left(x-\hat{x}\right)^{\top}\mathbf{w}_{V}$
@@
where the first inequality holds by Lemma I.3, the second inequality holds by the bounded assumption (Assumption 1), and the last inequality holds by the definition of $\delta=d/4-|\tilde{U}_{t}\cap V|$. ∎
@@ I.3 Technical Lemmas for Proposition 1
Lemma I.5 (Lemma 11 of Perivier and Goyal [42]).
@@
$\sum_{t=1}^{T}\sum_{i\in S_{t}}p_{t}(i|S_{t},\mathbf{w}^{\star})p_{t}(0|S_{t},\mathbf{w}^{\star})\leq\sum_{t=1}^{T}\kappa_{t}^{\star}+\textup{{Reg}}_{T}(\mathbf{w}^{\star}).$
@@
Then, we can rewrite the right-hand side of Equation (J.1) as follows:
@@
where $\bar{\mathbf{w}}_{t}$ is the convex combination of $\mathbf{w}^{\star}$ and $\mathbf{w}_{t+1}$. The first term of Equation (J.2) can be bound by
@@
Moreover, the second term of Equation (J.2) can be bounded as follows:
@@
where the first inequality holds by Lemma 1, and the last inequality holds by Lemma E.2. Combining Equation (J.3) and (J.4), we can bound the term $\sum_{t\in J_{1}}g_{t}(S_{t};\mathbf{w}^{\star})$ as follows:
@@
Now, by plugging Equation (J.5) and (J.6) into Equation (J.1), and using the fact that $\beta_{T}(\delta)=\tilde{\mathcal{O}}(\sqrt{d})$, we obtain that
@@
Furthermore, by applying the same analysis as in the proof of Theorem 2 in order to bound the second-order term in the Taylor expansion (term (B) in Equation (E.1)), we get
@@
For each instance, we sample the true parameter $\mathbf{w}^{\star}$ from a uniform distribution in $1/\sqrt{d},1/\sqrt{d}]^{d}$. For the context features $x_{ti}$, we sample each $x_{ti}$ independently and identically distributed (i.i.d.) from a multivariate Gaussian distribution $\mathcal{N}(\mathbf{0}_{d},\mathbf{I}_{d})$ and clip it to range $1/\sqrt{d},1/\sqrt{d}]^{d}$. Therefore, we ensure that $\|\mathbf{w}^{\star}\|_{2}\leq 1$ and $\|x_{ti}\|_{2}\leq 1$, satisfying Assumption 1. For each experimental configuration, we conducted $20$ independent runs for each instance and reported the average cumulative regret (Figure 1) and runtime per round (Figure K.1) for each algorithm. The error bars in Figure 1 and K.2 represent the standard deviations (1-sigma error). We have omitted the error bars in Figure K.1 because they are minimal.
@@
It appears that the authors may have intended to derive this equation using a first-order exact Taylor expansion. However, in MNL bandits, this equation generally does not hold. Consider a counterexample where $x_{ti}=0$, $x_{tj}\neq 0$ for $j\neq i$, and $\theta^{\star}\neq\theta_{t}$. Then, the left-hand side of Equation (L.1) is equals to 0 (since $x_{ti}=0$), but the right-hand side of Equation (L.1) is not 0, because the denominators of each $\mu_{i}(\mathbf{X}_{S_{t}}^{\top}\theta^{\star})$ and $\mu_{i}(\mathbf{X}_{S_{t}}^{\top}\theta_{t})$ differ. This equation only holds in special cases, such as when $K=1$, which corresponds to the logistic bandit case. Equation (16) in Agrawal et al. [4] serves as the foundation for the entire proof in their paper. Consequently, all subsequent results derived from it are also incorrect.